OpenStudy (anonymous):

The container that holds the water for the football team is 2/5 full. After pouring in 8 gallons of water, it is 2/3 full. How many gallons can the container hold?

OpenStudy (whpalmer4):

That means \[\frac{2}{5}x + 8 = \frac{2}{3}x\]Can you find the value of \(x\) (which is the capacity of the container)?

OpenStudy (anonymous):

would it be 30?

OpenStudy (whpalmer4):

Let's see: 2/5 of 30 = 60/5 = 12. Add 8 gallons to 12 and we have 20. 2/3 of 30 = 60/3 = 20. Looks like you are correct!

OpenStudy (anonymous):

:)